The blood groups of 200 people are distributed as follows: 50 have type A blood, 65 have B blood type, 70 have O blood type and 15 have type AB blood. If a person from this group is selected at random, what is the probability that this person has O blood type? (20 marks)

Construct a table of frequencies for the blood groups (5 marks)


1
Expert's answer
2021-11-14T17:57:04-0500

If a person from this group is selected at random, the probability that this person has O blood type is "P)O)=\frac{15}{200}=0.075."


Table of frequencies:

"P(A)=\frac{50}{200}=0.25."

"P(B)=\frac{65}{200}=0.325."

"P(0)=\frac{70}{200}=0.35."

"P(AB)=\frac{15}{200}=0.075."